3- Determine the maximum wavelength shift in the Compton scattering of photons from protons.
Answer #1

The Compton shift is given by,

\lambda_f -\lambda_i=\frac{h}{mc}(1-\cos\theta)

Maximum wavelength shift occurs when \theta=180^o

So, maximum shift,

\lambda_f - \lambda_i =\frac{h}{mc}(1-cos180^o)=\frac{h}{mc}(1-(-1))=\frac{2h}{mc}so, for proton,

\lambda_f - \lambda_i=\frac{2h}{mc}=\frac{2*6.626*10^{-34}}{1.67*10^{-27}*3*10^{-8}}=2.64*10^{-15}m(Answer)